OpenStudy (anonymous):

Help!!! You need a 30% alcohol solution. On hand, you have a 200 mL of a 15% alcohol mixture. You also have 80% alcohol mixture. How much of the 80% mixture will you need to add to obtain the desired solution?

OpenStudy (anonymous):

must be mixture day where is @FoolForMath when you need him?

OpenStudy (anonymous):

it is :(haha

OpenStudy (mertsj):

200 ml + x ml = (200+x) ml .15(200)+.80x=.30(200+x)

OpenStudy (mertsj):

Solve for x

OpenStudy (anonymous):

 \[\frac 3{10} \times 200 = 60 ml \]